# Asymptotically Unbiased Synthetic Control Methods by Density Matching

### v4

@@ 7.1 Conformal Inference on the Treatment Effect
We then apply our proposed method to this augmented dataset (comprising
$\left(Y_{0,1},\dots,Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$
and $(Y_{j,1})_{t\in\mathcal{T},j\in\mathcal{J}^{U}}$) to obtain adjusted weights $\widehat{\bm{w}}(\alpha)$ and an adjusted predictor of the counterfactual outcome, denoted $\widehat{Y}^{N}_{0,t}(\bm{\alpha})$.
@@
where $\mathcal{A}$ is a set of candidate values. Theorem 1 of Chernozhukov et al. (2021) guarantees (6).
7.2 Two-Sample Homogeneity Test
To test distributional treatment effects, consider the null and alternative hypotheses $H_{0}:p^{I}_{t}=p^{N}_{t}$ versus $H_{1}:p^{I}_{t}\neq p^{N}_{t}$. One can perform a two-sample homogeneity test in combination with our distributional treatment effect estimator. For instance, one may employ maximum mean discrepancy (MMD) (Gretton et al., 2012) to implement this hypothesis test.
@@ 8.1 Justification of Mixture Models
This result also implies that the mixture model (3) holds (Nazaret et al., 2023), aligning it with linear factor models. Hence, we obtain the following corollary for linear factor models similar to Ferman and Pinto (2021):
@@
Suppose that Assumptions 5.1, 6.1, and 8.18.4 hold. Then
$\bm{w}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\quad\text{as }T_{0}\to\infty,$
and $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ converges to
@@ 8.2 Auxiliary Covariates
^$\widehat{\ell}_{k}(\bm{w})=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(X_{0,t,k}-\sum_{j\in\mathcal{J}^{U}}w_{j}X_{j,t,k}\right).$
When these covariates are available, we can extend the DMSCM estimator as
$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0}}\coloneqq v_{0}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|+\sum_{k=1}^{K}v_{k}\widehat{\ell}^{2}_{k}(\bm{w}),$
where $\bm{v}\coloneqq(v_{k})_{k=0}^{K}$ are weights such that $v_{k}\geq 0$ and $\sum_{k=0}^{K}v_{k}=1$.
@@ 8.3 Distributional SCMs and Optimal Transport
8.4 Demeaned Density Matching SCM (D2MSCM)
Finally, we consider the case where the target distribution does not precisely satisfy (3), yet any deviation disappears after demeaning. Concretely, instead of (3), we assume
$p^{N}_{0,t}(y-c_{0})=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p^{N}_{j,t}(y-c_{j})\qquad\forall t\in\mathcal{T}.$
From the perspective of the linear factor model (1) and fine-grained models (Shi et al., 2022), this assumption permits $c_{j}\neq 0$ for some $j\in\mathcal{J}$. Under the previous mixture model, we implicitly required $c_{j}=0$ for all $j\in\mathcal{J}$, which may be restrictive in practice.
To estimate the SC weights in this setting, we first estimate $c_{j}$ by taking the average of $(Y_{j,t})_{t\in\mathcal{T}_{0}}$ and subtracting it from each $Y_{j,t}$. This demeaned approach is inspired by the proposal of Ferman and Pinto (2021) but is applied here to density matching. Let $h\in(0,1)$ be a fixed parameter. We estimate the weights as
$\widehat{\bm{w}}^{\mathrm{D2MSCM}}_{T_{0}}$ $\coloneqq\left(\widehat{w}^{\mathrm{D2MSCM}}_{j,T_{0}}\right)_{j\in\mathcal{J}^{U}}$
$\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\biggl{|}\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}-\overline{Y}^{N}_{0}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)^{\gamma}\right)\biggr{|}.$
We call this procedure demeaned DMSCM (D2SCM). Using these weights, we can estimate the treatment effect as
$\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\biggl{(}\overline{Y}^{N}_{0}+\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{D2MSCM}}_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)\biggr{)}.$
Figure 2: Results of simulation studies.
@@ 9 Experiments
We conduct experiments using both simulated and real-world data. Additional results can be found in Appendix 9.1.
@@ 9.1 Simulation Studies
For the treated unit, we let $Y^{N}_{0,t}\sim p_{0,t}(\bm{z})$ for $t\in[t_{0},T_{1}]$ and set $Y^{I}_{0,t}=20+Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$, implying $\tau_{0,t}=20$. We draw $\bm{w}^{*}$ from a uniform distribution over $[0,1]^{J}$ and then normalize it so that its entries sum to $1$.
We apply DMSCM and the method of Abadie and Gardeazabal (2003) (referred to here as Abadie) $1000$ times and evaluate the estimation error of the treatment effect. We plot the estimation error box plots in Figure 3, indicate that our proposed method achieves smaller estimation errors. Moreover, the estimation error declines as $G$ increases.
Next, we consider another set of simulation studies under conditions similar to those above, but with the following changes: $J\in\{1,5,10,15,20,\dots,35,40,45,50\}$, $G\in\{2,3,5,10\}$, $T_{0}=30$, $T_{1}=1000$, and $K=5$. All other settings remain as in Section 9.1.
We again apply DMSCM and Abadie $1000$ times and compute the error of the treatment effect estimation (Figure 3). We focus on how the mean squared estimation error changes as $J$ grows. While the estimation error under the method of Abadie (2002) grows with $J$, our proposed method does not exhibit such growth. Additionally, our method maintains lower estimation errors than the method of Abadie (2002).
A similar pattern appears in distributional treatment effect estimation. We measure the discrepancy between the true and estimated distributions using the Maximum Mean Discrepancy (MMD, Gretton et al., 2012), a probability metric based on kernel embeddings. In the top panel of Figure 4, we show how MMD varies with $J$. As before, our proposed methods outperform the method of Abadie (2002) in this setting. In the bottom panel of Figure 4, we plot MMD for $G\in\{2,10,50,100\}$. As $G$ grows (i.e., as more moment conditions are imposed), the predictive performance improves.
@@
Table 2: Results of treatment effect estimation and conformal inference. Estimated treatment effect $\widehat{\tau}_{0,t}$ and $90\%$ confidence interval $(\underline{\tau}_{0,t},\overline{\tau}_{0,t})$.
Terrorist conflict in Basque country.
Abadie DMSCM
$\gamma$- $2$ $5$ $10$ $100$
$\widehat{\tau}_{0,t}$ $-0.800$ $-0.845$ $-1.054$ $-0.279$ $-0.138$
$\overline{\tau}_{0,t}$ $0.522$ $0.699$ $0.495$ $0.521$ $0.578$
$\underline{\tau}_{0,t}$ $-2.587$ $-2.150$ $-2.605$ $-2.616$ $-2.590$
Tobacco control program in California.
Abadie DMSCM
$\gamma$- $2$ $5$ $10$ $100$
$\widehat{\tau}_{0,t}$ $-26.148$ $-27.744$ $-32.306$ $-35.080$ $-36.879$
$\overline{\tau}_{0,t}$ $8.698$ $7.852$ $8.898$ $9.544$ $11.748$
$\underline{\tau}_{0,t}$ $-36.798$ $-29.268$ $-34.783$ $-37.308$ $-35.243$
Reunification of Germany.
Abadie DMSCM
$\gamma$- $2$ $5$ $10$ $100$
$\widehat{\tau}_{0,t}$ $-1492.374$ $-1253.616$ $-1419.120$ $-1324.448$ $-510.765$
$\overline{\tau}_{0,t}$ $2635.583$ $-1008.822$ $-591.420$ $-716.376$ $2327.944$
$\underline{\tau}_{0,t}$ $-2390.413$ $-2738.232$ $-1989.323$ $-2018.878$ $-2625.129$
@@ 9.2 Empirical Analysis
We now present empirical studies based on Abadie and Gardeazabal (2003), Abadie et al. (2010), and Abadie et al. (2015).
Terrorist conflict in the Basque Country. Abadie and Gardeazabal (2003) investigate the effect of terrorist conflict in the Basque Country on gross domestic product (GDP). In this example, $J=16$, $K=14$, and $(t_{0},T_{0},T_{1})=(1955,1970,1997)$. Basque Euskadi Ta Askatasuna (Basque Homeland and Liberty; ETA) was founded in 1959 as a Basque separatist organization in Spain, seeking an independent Basque state primarily through terrorist activities in the Basque Country region. Abadie and Gardeazabal (2003) use annual Spanish regional-level panel data from 1955 to 1997, with a particular focus on the 1970s when ETA’s terrorist activity was significant. The dataset includes fifteen years of pre-intervention observations. The set of untreated units consists of $17$ other Spanish regions. Each region’s record contains the outcome of interest (GDP per capita) and additional covariates (investment rate, population density, five sectoral output measures, and four measures of human capital), all averaged from 1955 to 1968.
Tobacco control program in California. Abadie et al. (2010) study the effect of a large tobacco control program adopted in California in 1988 (described in the Introduction). Here, $J=20$, $K=5$, and $(t_{0},T_{0},T_{1})=(1970,1989,2000)$.
@@
Separately, Abadie et al. (2010) also investigate the causal effect of Proposition 99, a comprehensive tobacco control program adopted by California. They use annual state-level panel data from 1970 to 2000. The program was implemented in January 1989, providing eighteen years of pre-intervention observations. Four states that adopted large-scale tobacco control programs (Massachusetts, Arizona, Oregon, and Florida) and states that raised cigarette taxes by at least \$0.50 between 1989 and 2000 (Alaska, Hawaii, Maryland, Michigan, New Jersey, New York, and Washington) as well as the District of Columbia are excluded from the set of 38 control states. The outcome of interest is annual per capita cigarette sales in packs; covariates include the average retail price of cigarettes, log of state income per capita, the percentage of the population aged 15–24, and per capita beer consumption, all averaged between 1980 and 1988.
Results. Table 2 reports the estimated treatment effects and 90% confidence intervals. Figures 6 and 6 show the estimated treatment effects and the distributional treatment effects, respectively.
In each dataset, we plot the counterfactual outcomes and distributions in Figures 6 and 6. We also perform conformal inference on the sharp null hypothesis $H_{0}:\tau_{0,(T_{0}+1)}=\cdots=\tau_{0,T_{1}}=0$ for the treatment effect. For both Abadie’s method and DMSCM, we present the 90% confidence interval of the treatment effect in Table 2. Our proposed method fits the observed outcomes well in the pre-treatment period ($t\in\mathcal{T}_{0}$) and, accordingly, produces plausible counterfactual outcomes and distributions. In the conformal inference analysis, our method does not reject the null hypothesis in the same manner as Abadie’s estimator, except for the DMSCM in the California tobacco control example.

### v5

@@ 5.6 Inference
We then apply our proposed method to this augmented dataset, comprising $\left(Y_{0,1},\dots,Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$ and $(Y_{j,1})_{t\in\mathcal{T},j\in\mathcal{J}^{U}}$, to obtain adjusted weights $\widehat{\bm{w}}(\alpha)$ and an adjusted predictor of the counterfactual outcome, denoted $\widehat{Y}^{N}_{0,t}(\bm{\alpha})$.
@@
where $\mathcal{A}$ is a set of candidate values. Theorem 1 of Chernozhukov et al. (2021) guarantees (6). We summarize the procedure in Algorithm 2
Remark 5.11 (Conformal inference and randomization inference).
Conformal inference provides confidence intervals under a stationarity assumption and a prediction model for the counterfactual outcome. In contrast, randomization inference requires random assignment and prediction models for post treatment outcomes of all units. These differences are recognized advantages of conformal inference (Chernozhukov et al., 2021), and the method has recently been widely used in SCMs (Ben-Michael et al., 2021). Furthermore, conformal inference can itself be interpreted as a variant of randomization inference (Ritzwoller et al., 2025).
Algorithm 2 Conformal prediction.
Sharp null hypothesis: $H_{0}:\tau_{0,t}=\alpha_{t}$ for some $t\in\mathcal{T}_{t}$ and $\bm{\alpha}=(\alpha_{t})_{t\in\mathcal{T}_{1}}$.
 Make an augmented dataset $\left(Y_{0,1},\dots,Y_{0,T_{0}},Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$.
 We apply the MMSCM by regarding the datasets $\left(Y_{0,1},\dots,Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$ and $(Y_{j,1})_{t\in\mathcal{T},j\in\mathcal{J}^{U}}$ as the pretreatment dataset.
 Using the estimated weight, we obtain estimate of the counterfactual outcome $\widehat{Y}^{N}_{0,t}(\bm{\alpha})$.
 Obtain $u_{t}(\bm{\alpha})=Y_{0,t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{0}$, and $u_{t}(\bm{\alpha})=Y^{I}_{0,t}-\alpha_{t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{1}$.
 Compute the $p$-value as $\widehat{p}(\bm{\alpha})=1-\widehat{F}(\bm{\alpha})$ and confidence interval as $\widehat{C}^{\mathrm{conf}}_{1-\xi}=\left\{\bm{\alpha}\in\mathcal{A}\bigm|\widehat{p}(\bm{\alpha})>\xi\right\}$.
5.7 Distributional Treatment Effect Estimation
To estimate the distributional treatment effect, we approximate the counterfactual distribution $p_{0,t}(y)$ using a bootstrap-like simulation-based method. Specifically, we resample $L>1$ observations $\{Y^{l,\dagger}_{0}\}_{l=1}^{L}$ from $\{Y_{0,t},Y_{1,t},\dots,Y_{J,t}\}_{t\in\mathcal{T}_{1}}$ as follows: for each $l\in\mathcal{L}\coloneqq\{1,2,\dots,L\}$,
  1. 1.
  2. 2.
[…]
The empirical distribution of $\{Y^{l,\dagger}_{0}\}_{l=1}^{L}$ serves as an estimator of the distribution of $\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}Y_{0,t}$, as shown in the following theorem. The proof is provided in Appendix B. Thus, when our primary focus is on $\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p_{0,t}(y)$, the above procedure enables us to achieve this goal.
Theorem 5.12.
Suppose that the assumptions in Theorem 5.8 hold. Define
  • $G_{T_{0},L}\coloneqq\frac{1}{L}\sum^{L}_{l=1}\mathbbm{1}[Y^{l,\dagger}_{0}\leq y]$.
  • $F_{Y^{N}_{j,t},T_{0}}(y)\coloneqq\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\mathbbm{1}[Y^{N}_{j,t}\leq y]$.
  • $\widehat{F}_{Y^{N}_{0,t},T_{0}}(y)\coloneqq\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{MMSCM}}_{j,T_{0},G}F_{Y^{N}_{j,t},T_{0}}(y)$.
Then, the following holds:
$\sup_{y}\big|G_{T_{0},L}(y)-F_{Y^{N}_{0,t}}(y)\big|\xrightarrow{\mathrm{p}}0\quad(L\to\infty,T_{0}\to\infty),$
where we recall that $F_{Y^{N}_{0,t}}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F_{Y^{N}_{j,t}}(y)\quad\forall y\in\mathbb{R}$.
Two-sample homogeneity test. To test distributional treatment effects, consider the null and alternative hypotheses $H_{0}:p^{I}_{t}=p^{N}_{t}$ versus $H_{1}:p^{I}_{t}\neq p^{N}_{t}$. One can perform a two-sample homogeneity test in combination with our distributional treatment effect estimator. For instance, one may employ maximum mean discrepancy (MMD) (Gretton et al., 2012) to implement this hypothesis test.
@@ 6.1 Justification of Mixture Models
This result also implies that the mixture model (3) holds (Nazaret et al., 2023), aligning it with linear factor models. Hence, we obtain the following corollary for linear factor models similar to Ferman & Pinto (2021):
@@
Suppose that Assumptions 5.4, 5.7, and 6.16.4 hold. Then
$\bm{w}^{\mathrm{MMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\quad\text{as }T_{0}\to\infty,$
and $\widehat{\tau}^{\mathrm{MMSCM}}_{0,t}$ converges to
@@ 6.2 Auxiliary Covariates
$\widehat{\ell}_{k}(\bm{w})=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(X_{0,t,k}-\sum_{j\in\mathcal{J}^{U}}w_{j}X_{j,t,k}\right).$
When these covariates are available, we can extend the MMSCM estimator as
$\widehat{\bm{w}}^{\mathrm{MMSCM}}_{T_{0}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}b_{0}\sum_{\gamma=1}^{G}v_{\gamma}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|+\sum_{k=1}^{K}b_{k}\widehat{\ell}^{2}_{k}(\bm{w}),$
where $(b_{k})_{k=0}^{K}$ are weights such that $b_{k}\geq 0$ and $\sum_{k=0}^{K}b_{k}=1$.
@@ 6.3 Distributional SCMs and Optimal Transport
Figure 2: Results of simulation studies. Each result represents estimation error ($\tau_{0,t}-\widehat{\tau}$ for an estimator $\widehat{\tau}$). The orange horizontal line denotes the median, and the triangle mark denotes the mean.
@@ 7 Experiments
We conduct experiments using both simulated and real-world data. In all experiments, we set $v_{\gamma}=1$ for the MMSCM.
@@ 7.1 Simulation Studies
For the treated unit, we let $Y^{N}_{0,t}\sim p_{0,t}(\bm{z})$ for $t\in[t_{0},T_{1}]$ and set $Y^{I}_{0,t}=20+Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$, implying $\tau_{0,t}=20$. We draw $\bm{w}^{*}$ from a uniform distribution over $0,1]^{J}$ and then normalize it so that its entries sum to $1$.
We apply MMSCM, the method proposed in Abadie et al. (2010) (referred to here as Abadie), and DiSCo in Gunsilius (2023). We iterate trials $100$ times and evaluate the estimation error of the treatment effect. We plot the estimation error box plots in Figure 2 with medians and means, indicate that our proposed method achieves smaller estimation errors. Moreover, the estimation error declines as $G$ increases.
Next, we consider another set of simulation studies under conditions similar to those above, but with the following changes: $J\in\{1,5,10,15,20,25,30\}$, $G\in\{2,5,10\}$, $T_{0}=30$, $T_{1}=1000$, and $K=5$. All other settings remain as in Section 7.1.
We again apply MMSCM and Abadie $100$ times and compute the error of the treatment effect estimation (Figure 3). We focus on how the mean squared estimation error changes as $J$ grows. While the estimation error under the method of Abadie (2002) grows with $J$, our proposed method does not exhibit such growth. Additionally, our method maintains lower estimation errors than the method of Abadie (2002).
A similar pattern appears in distributional treatment effect estimation. We measure the discrepancy between the true and estimated distributions using the Maximum Mean Discrepancy (MMD, Gretton et al., 2012), a probability metric based on kernel embeddings. In the top panel of Figure 4, we show how MMD varies with $J\in\{1,5,10,15,20,\dots,35,40,45\}$. As before, our proposed methods outperform the method of Abadie (2002) in this setting. In the bottom panel of Figure 4, we plot MMD for $G\in\{2,10,50,100\}$. As $G$ grows (i.e., as more moment conditions are imposed), the predictive performance improves. Note that in this experiment, we included the results of Abadie et al. (2010) for reference only, since it is not a method for distributional SCM. Our focus is on the gains induced by choosing a larger $G$.
@@ 7.2 Empirical Analysis
We now present empirical studies based on the datasets and settings in Abadie & Gardeazabal (2003), Abadie et al. (2010), and Abadie et al. (2015). In this analysis, we also compare our method with the prediction intervals for SCMs proposed by Cattaneo et al. (2021), which provide a robust uncertainty evaluation for SCMs. Although this prediction interval cannot be applied to our method due to differences in the underlying models, it serves as a promising alternative to conformal prediction. Therefore, we include its results for reference. Note that while we used our own implementation for our method, Abadie’s method, and DiSCo, we used the pcsi library for the prediction intervals, as published by Cattaneo et al. (2022).
Terrorist conflict in the Basque Country. Abadie & Gardeazabal (2003) investigate the effect of terrorist conflict in the Basque Country on gross domestic product (GDP). In this example, $J=16$, $K=14$, and $(t_{0},T_{0},T_{1})=(1955,1970,1997)$. Basque Euskadi Ta Askatasuna (Basque Homeland and Liberty; ETA) was founded in 1959 as a Basque separatist organization in Spain, seeking an independent Basque state primarily through terrorist activities in the Basque Country region. Abadie & Gardeazabal (2003) use annual Spanish regional-level panel data from 1955 to 1997, with a particular focus on the 1970s when ETA’s terrorist activity was significant. The dataset includes fifteen years of pre-intervention observations. The set of untreated units consists of $17$ other Spanish regions. Each region’s record contains the outcome of interest (GDP per capita) and additional covariates (investment rate, population density, five sectoral output measures, and four measures of human capital), all averaged from 1955 to 1968.
Tobacco control program in California. Abadie et al. (2010) also investigate the causal effect of Proposition 99, a comprehensive tobacco control program adopted by California. They use annual state-level panel data from 1970 to 2000. The program was implemented in January 1989, providing eighteen years of pre-intervention observations. Four states that adopted large-scale tobacco control programs (Massachusetts, Arizona, Oregon, and Florida) and states that raised cigarette taxes by at least \$0.50 between 1989 and 2000 (Alaska, Hawaii, Maryland, Michigan, New Jersey, New York, and Washington) as well as the District of Columbia are excluded from the set of 38 control states. The outcome of interest is annual per capita cigarette sales in packs; covariates include the average retail price of cigarettes, log of state income per capita, the percentage of the population aged 15–24, and per capita beer consumption, all averaged between 1980 and 1988. Here, $J=20$, $K=5$, and $(t_{0},T_{0},T_{1})=(1970,1989,2000)$.
@@
Results. For each dataset, Figures 5 and 6 show the estimated treatment effects and the distributional treatment effects, respectively. In Appendix C, Tables 24 report the estimated treatment effects and 90% confidence bands obtained via conformal inference. We also plot the confidence bands in Figures 79.
Our proposed method fits the observed outcomes well in the pre-treatment period ($t\in\mathcal{T}_{0}$) and, accordingly, produces plausible counterfactual outcomes and distributions.
In the conformal inference analysis, when we set the sharp null hypothesis as $H_{0}:\tau_{0,(T_{0}+1)}=\cdots=\tau_{0,T_{1}}=0$ for the treatment effect, none of the methods reject the null hypothesis, except for the “California tobacco control” dataset. Prediction intervals proposed in Cattaneo et al. (2021) shows tight intervals for every settings. Note that we cannot directly compare the performance of their method and ours due to differences in the underlying models.